OpenStudy (anonymous):

I'm in 7th grade and doing POPS high school Algebra I. Can someone help me?

OpenStudy (anonymous):

sure post the problem

OpenStudy (anonymous):

post the problem !

OpenStudy (lukecrayonz):

Post the problem and I can help.

OpenStudy (anonymous):

-6x+7=-5 and the answer is 2 how do I show the work?

jhonyy9 (jhonyy9):

subtract from both sides 7 -6x+7-7=-5-7 -6x=-12 now divide both sides by -6 -6x/-6 = -12/-6 x=2

OpenStudy (anonymous):